Any help on this would be appreciated. Pleasse show your work. Thanks
Let F(u)=u^2-u-1 Find
(A)F(10) (B)F(u^2) (C)(5u) (D)5F(u)
Thanks once again.

Answer by Alan3354(69443)

Let F(u)=u^2-u-1 Find
(A)F(10)
Wherever there's a u, sub 10
F(10) = 10^2 - 10 - 1 = 89
--------
Do the same for the others.
(B)F(u^2) (C)(5u) (D)5F(u)